OpenStudy (anonymous):

Find two consecutive even numbers such that the difference of one-half the larger number and two-fifths the smaller number is equal to five. 38 and 40 40 and 42 42 and 44

OpenStudy (raden):

let that numbers are x and (x+2) respectively. so, 1/2 (x+2) - 2/5 (x) = 5 multiply 10 on both sides, get 10 * 1/2 (x+2) - 10 * 2/5 (x) = 10 * 5 5(x+2) - 4x = 50 5x + 10 - 4x = 50 x + 10 = 50 x = 50 -10 x = 40 and the 2nd number = x+2 = 40 + 2 = 42 they are 40 and 42

OpenStudy (anonymous):

thank you